# SNE-RoadSegV2: Advancing Heterogeneous Feature Fusion and Fallibility Awareness for Freespace Detection

### v1

@@ Abstract
To address the aforementioned limitations, we first dive deeper into the discriminative feature fusion strategies presented in recent universal semantic segmentation studies [1, 7]. Subsequently, we introduce a novel heterogeneous feature fusion block (HFB) to process the RGB and surface normal features, which are encoded using two independent Swin Transformer backbones [10]. Our technical contributions in this part are three-fold: (1) a holistic attention module (HAM) to model the interdependencies between heterogeneous features across three dimensions (spatial, channel, and scale); (2) a heterogeneous feature contrast descriptor (HFCD) to effectively underscore both the shared and unique characteristics within the holistically-attentive features; (3) an affinity-weighted feature recalibrator (AWFR) to jointly emphasize and suppress heterogeneous features prior to their input into the decoder. Additionally, we contribute to a lightweight yet more effective decoder, which incorporates inter-scale skip connections while pruning redundant ones. Our decoder demonstrates greater feature decoding capabilities, while simultaneously reducing computational complexity. Finally, we design two new loss functions based on semantic annotations and depth data to provide deeper supervision during our model training process. This contribution also results in improved overall performance, particularly in error-prone areas. The effectiveness of each contribution is validated through extensive experiments conducted across public datasets. Excitingly, our proposed freespace detection framework SNE-RoadSegV2, as shown in Fig. 1, which incorporates all these innovative components, wins the 1st place on the KITTI Road leaderboard [11]. In a nutshell, our contributions are as follows:
@@
We introduce HFB, consisting of an HAM, an HFCD, and an AWFR, for comprehensive heterogeneous feature description, recalibration, and fusion, resulting in more coherent feature representations.
@@ III-A Architecture Overview
Fig. 1 provides readers with an overview of the SNE-RoadSegV2 architecture, consisting of three key elements: duplex feature embedding, heterogeneous feature fusion, and lightweight yet effective feature decoding. A pair of input RGB image and surface normal map , translated from a depth image using a surface normal estimator (SNE) [1], are first tokenized into non-overlapping patches and transformed into a high-dimensional feature space through a trainable linear projection in the patch embedding module [10]. The embedded features are subsequently fed into the Swin Transformer blocks [10] to produce hierarchical heterogeneous features and . Each pair of heterogeneous features undergoes a comprehensive fusion process through HFB. Finally, a lightweight yet more effective decoder which incorporates both inter-scale and intra-scale skip connections is designed to further boost the efficiency and accuracy of freespace detection. The proposed architecture is trained by minimizing a loss with fallibility-awareness incorporated. The following subsections will provide a detailed explanation of the HFB, decoder, and loss functions in sequence.
@@ III-B Heterogeneous Feature Fusion Block
The core problem of feature encoding for freespace detection revolves around the effective fusion of heterogeneous features extracted from various data sources. As detailed in Sec. I, prior investigations, e.g. , [1, 7, 23, 21] fused these heterogeneous features without appropriate discrimination between their inherent differences. Our HFB is, thus, designed to overcome this limitation. As depicted in Fig. 2, meaningful heterogeneous features are first selectively emphasized and suppressed across three dimensions (spatial, channel, and scale) through an HAM. These features are further enhanced through an HFCD, which improves the representations of their shared and distinct characteristics. The original heterogeneous features are ultimately weighted through an AWFR to emphasize the aspects that are important to both or either of the features.
@@ III-B3 Affinity-Weighted Feature Recalibrator
Constructing a volume that contains element-wise weights to jointly recalibrate (emphasize and de-emphasize) heterogeneous features is another significant contribution in our designed HFB. As and describe the contrasting aspects between the heterogeneous features at the spatial and scale-channel levels, respectively, we employ these two volumes to construct an affinity volume as follows:
@@
where , , and denote convolutional layers. Compared to other SoTA heterogeneous feature fusion strategies, our proposed HFB adaptively assigns weights to the original features, taking into account both the elements of agreement and disagreement in importance as determined by HFCD and AWFR. Such a way of seeking common ground and while preserving differences enhances the comprehensiveness of feature fusion for freespace detection. The superior performance of HFB and the effectiveness of its each component are demonstrated in Sec. IV-D.
@@ IV-C Comparison with State-of-the-art Methods
The quantitative comparisons among SoTA freespace detection algorithms on the KITTI Semantics [11] and VKITTI2 datasets are presented in Tables III and IV, respectively, while their qualitative comparisons on these two datasets are provided in Figs. 7 and 8. These results suggest that our proposed SNE-RoadSegV2 outperforms all other SoTA approaches, with an increase in Fsc by 1.24-4.01% on the KITTI Semantics dataset, and by 1.08-4.08% on the VKITTI2 dataset. It is noteworthy that our SNE-RoadSegV2 demonstrates greater robustness, particularly near/on semantic transition regions or in areas with surface normals similar to that of the road surface. The achieved improvements are primarily attributed to our heterogeneous feature fusion strategy and fallibility-aware loss functions.
@@ IV-D2 Encoder
We first explore the rationality of each component in HFB. As presented in Table VIII, we sequentially remove each component from HFB to quantify its impact on the overall performance. It is evident that each component in HFB contributes to an improvement in the overall performance, and the network achieves its peak performance when all components (HAM, HFCD, and AWFR) are integrated into HFB, which demonstrates the effectiveness of our design.
Additionally, we compare HFB with other SoTA heterogeneous feature fusion strategies. As shown in Table IX, HFB outperforms other compared methods on both datasets, with improvements of up to 1.26% in Fsc and 1.67% in IoU, respectively. These compelling results can be attributed to our novel contributions, particularly the exploitation of both shared and distinct characteristics of heterogeneous features in HFCD, and the effective feature recalibration based on the affinity volume constructed in AWFR. Another qualitative comparison illustrated in Fig. 6 further validates the effectiveness of HFB. Obviously, HFB produces more coherent feature representations with clearer segmentation boundaries, compared to the indiscriminate feature fusion strategy adopted in SNE-RoadSeg.

### v2

@@ Abstract
To address the aforementioned limitations, we first dive deeper into the discriminative feature fusion strategies presented in recent universal semantic segmentation studies [1, 7]. Subsequently, we introduce a novel heterogeneous feature fusion block (HF2B) to process the RGB and surface normal features, which are encoded using two independent Swin Transformer backbones [10]. Our technical contributions in this part are three-fold: (1) a holistic attention module (HAM) to model the interdependencies between heterogeneous features across three dimensions (spatial, channel, and scale); (2) a heterogeneous feature contrast descriptor (HFCD) to effectively underscore both the shared and unique characteristics within the holistically-attentive features; (3) an affinity-weighted feature recalibrator (AWFR) to jointly emphasize and suppress heterogeneous features prior to their input into the decoder. Additionally, we contribute to a lightweight yet more effective decoder, which incorporates inter-scale skip connections while pruning redundant ones. Our decoder demonstrates greater feature decoding capabilities, while simultaneously reducing computational complexity. Finally, we design two new loss functions based on semantic annotations and depth data to provide deeper supervision during our model training process. This contribution also results in improved overall performance, particularly in error-prone areas. The effectiveness of each contribution is validated through extensive experiments conducted across public datasets. Excitingly, our proposed freespace detection framework SNE-RoadSegV2, as shown in Fig. 1, which incorporates all these innovative components, wins the 1st place on the KITTI Road leaderboard [11]. In a nutshell, our contributions are as follows:
@@
We introduce HF2B, consisting of an HAM, an HFCD, and an AWFR, for comprehensive heterogeneous feature description, recalibration, and fusion, resulting in more coherent feature representations.
@@ III-A Architecture Overview
Fig. 1 provides readers with an overview of the SNE-RoadSegV2 architecture, consisting of three key elements: duplex feature embedding, heterogeneous feature fusion, and lightweight yet effective feature decoding. A pair of input RGB image and surface normal map , translated from a depth image using a surface normal estimator (SNE) [1], are first tokenized into non-overlapping patches and transformed into a high-dimensional feature space through a trainable linear projection in the patch embedding module [10]. The embedded features are subsequently fed into the Swin Transformer blocks [10] to produce hierarchical heterogeneous features and . Each pair of heterogeneous features undergoes a comprehensive fusion process through HF2B. Finally, a lightweight yet more effective decoder which incorporates both inter-scale and intra-scale skip connections is designed to further boost the efficiency and accuracy of freespace detection. The proposed architecture is trained by minimizing a loss with fallibility-awareness incorporated. The following subsections will provide a detailed explanation of the HF2B, decoder, and loss functions in sequence.
@@ III-B Heterogeneous Feature Fusion Block
The core problem of feature encoding for freespace detection revolves around the effective fusion of heterogeneous features extracted from various data sources. As detailed in Sec. I, prior investigations, e.g. , [1, 7, 23, 21] fused these heterogeneous features without appropriate discrimination between their inherent differences. Our HF2B is, thus, designed to overcome this limitation. As depicted in Fig. 2, meaningful heterogeneous features are first selectively emphasized and suppressed across three dimensions (spatial, channel, and scale) through an HAM. These features are further enhanced through an HFCD, which improves the representations of their shared and distinct characteristics. The original heterogeneous features are ultimately weighted through an AWFR to emphasize the aspects that are important to both or either of the features.
@@ III-B3 Affinity-Weighted Feature Recalibrator
Constructing a volume that contains element-wise weights to jointly recalibrate (emphasize and de-emphasize) heterogeneous features is another significant contribution in our designed HF2B. As and describe the contrasting aspects between the heterogeneous features at the spatial and scale-channel levels, respectively, we employ these two volumes to construct an affinity volume as follows:
@@
where , , and denote convolutional layers. Compared to other SoTA heterogeneous feature fusion strategies, our proposed HF2B adaptively assigns weights to the original features, taking into account both the elements of agreement and disagreement in importance as determined by HFCD and AWFR. Such a way of seeking common ground and while preserving differences enhances the comprehensiveness of feature fusion for freespace detection. The superior performance of HF2B and the effectiveness of its each component are demonstrated in Sec. IV-D.
@@ IV-C Comparison with State-of-the-art Methods
The quantitative comparisons among SoTA freespace detection algorithms on the KITTI Semantics [11] and VKITTI2 datasets are presented in Tables III and IV, respectively, while their qualitative comparisons on these two datasets are provided in Figs. 9 and 10. These results suggest that our proposed SNE-RoadSegV2 outperforms all other SoTA approaches, with an increase in Fsc by 1.24-4.01% on the KITTI Semantics dataset, and by 1.08-4.08% on the VKITTI2 dataset. It is noteworthy that our SNE-RoadSegV2 demonstrates greater robustness, particularly near/on semantic transition regions or in areas with surface normals similar to that of the road surface. The achieved improvements are primarily attributed to our heterogeneous feature fusion strategy and fallibility-aware loss functions.
@@ IV-D2 Encoder
We first explore the rationality of each component in HF2B. As presented in Table VIII, we sequentially remove each component from HF2B to quantify its impact on the overall performance. It is evident that each component in HF2B contributes to an improvement in the overall performance, and the network achieves its peak performance when all components (HAM, HFCD, and AWFR) are integrated into HF2B, which demonstrates the effectiveness of our design.
Additionally, we compare HF2B with other SoTA heterogeneous feature fusion strategies. As shown in Table IX, HF2B outperforms other compared methods on both datasets, with improvements of up to 1.26% in Fsc and 1.67% in IoU, respectively. These compelling results can be attributed to our novel contributions, particularly the exploitation of both shared and distinct characteristics of heterogeneous features in HFCD, and the effective feature recalibration based on the affinity volume constructed in AWFR. Another qualitative comparison illustrated in Fig. 6 further validates the effectiveness of HF2B. Obviously, HF2B produces more coherent feature representations with clearer segmentation boundaries, compared to the indiscriminate feature fusion strategy adopted in SNE-RoadSeg.